Can you help me Please. Find R if a=7, b=3, c=2, d=4 and e=5: R = a + 2b + 5c + d + 2e. The book says the answer is R = 33. I get R = 37. Here is how I solved the problem.
7 (1) + 3 (2) + 2 (5) + 4 + 5 (2) simplify that
7 + 6 + 10 + 4 + 10. Multiply R = 37
The book says should be R = 33. What am I doing wrong. Is the book have a typo error. If this works out im looking for a tutor with math. Thank you for your time.

Found 2 solutions by ewatrrr, richwmiller:
Answer by ewatrrr(24785)

R = a + 2b + 5c + d + 2e
7 (1) + 3 (2) + 2 (5) + 4 + 5 (2)
R = 7 + 6 + 10 + 4 + 10 = 37

Answer by richwmiller(17219)